# Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation

### v4

@@ Abstract
Recently, Large Language Models (LLMs) have shown impressive performance in various character understanding tasks. However, the extensive pre-training corpora used by LLMs raise concerns that they may rely on memorizing popular fictional works rather than genuinely understanding and reasoning about them. Prior work has addressed memorization in LLM reasoning tasks, but these approaches do not directly apply to character understanding. To address this, we leverage the cognitive science concepts of "gist memory" and "verbatim memory", where verbatim memory corresponds to exact recall, while gist memory captures underlying meaning. We design a hard-soft intervention framework, where the hard setting disrupts memorization by perturbing key entities, while the soft setting maintains the original text but guides the model toward reasoning. Our experiments across six character understanding tasks with GPT-4o, LLaMA3.3-70B, and DeepSeek V3 demonstrate that reducing verbatim memory reliance shifts model behavior toward reasoning. Our method leads to a performance drop of up to 45.8% when models can no longer depend on memorization, highlighting the extent to which existing benchmarks inflate performance through data contamination. Additionally, we show that prompting techniques can systematically alter a model’s response strategy, making it more dependent on either memorization or reasoning.
@@ 1 Introduction
Character understanding requires the models to grasp the characters’ personalities and traits while reading a story, which aims to evaluate the models’ capability to comprehend personalities, goals, roles, and relationships of various fictional characters Yu et al. (2023a); Vallurupalli et al. (2024); Shen et al. (2024); Brahman et al. (2021). However, the impressive performance of LLMs on such benchmarks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023) raises the question: does it reflect genuine understanding or merely strong memorization Chang et al. (2023a)? For example, in the TV show speaker guessing task shown in Figure  1, the model is expected to identify Sheldon based on its memory of the character from previous episodes (e.g., preparing lessons for Howard). However, if the model simply memorizes the exact words, it could make predictions based on a very short sentence by directly extracting the speaker information from it.
To truly stimulate a model’s character understanding ability, it is crucial to understand how the model may be generating its responses. This can be especially challenging given the extensive and opaque pre-training corpora used by LLMs, making it particularly severe for character understanding tasks, as many benchmarks are derived from well-known and widely consumed fiction or scripts Sang et al. (2022a). We hypothesize that this can lead to a form of memorization called “verbatim memory,” which encodes precise details and remembers every word exactly Huang et al. (2024); Ippolito et al. (2022)—rather than reasoning over recalled overarching patterns (what is referred to as “gist memory” Reyna and Brainerd (1998)).
Figure 1: This figure illustrates the distinct mechanisms of gist and verbatim memory processes in the speaker-guessing task for a script. Verbatim memory (bottom right) aims to recall as much specific information as possible, including who said each sentence. If the memorization is not precise enough, the model fails to provide the correct answer. In contrast, gist memory (top right) focuses on extracting core information, such as character relationships, personalities, and key events. By relying on gist memory, the model is encouraged to engage in reasoning rather than purely relying on memorization.
Motivated by the concept of gist memory, we propose a prompt-based approach to mitigate model reliance on verbatim memorization in character understanding tasks. We evaluate our method on GPT-4o, LLaMA3-70B, and DeepSeek V3, introducing a hard-soft setting framework: the hard setting blocks memorization by replacing character names, while the soft setting retains the original text but employs gist-based prompts to encourage reasoning over recall. Our ablation study confirms the effectiveness of both interventions.
Building on these findings, we extend our approach to five character understanding tasks, designing task-specific gist prompts and evaluating models under both settings. Our results indicate that the Gist+NR approach best reduces memorization dependency.
In summary, we outline our study and key contributions as follows:
@@
    Character understanding tasks heavily rely on verbatim memorization. Across six benchmark tasks, all evaluated LLMs show significant performance drops when explicit memorization is disrupted, confirming that these tasks often assess recall rather than true reasoning (Table 6.2).
@@
    Gist-based prompting effectively reduces memorization dependency. Introducing gist-based prompts systematically lowers reliance on verbatim recall across all tasks, encouraging models to use reasoning instead of mechanical memorization (Section 6).
@@
    Verbatim and gist-based prompting modulate model response strategies. Models adjust their response strategies based on the prompting style—verbatim-based prompts reinforce reliance on memorization, while gist-based prompts encourage reasoning-driven answers (Section 4.3).
@@ 2 Related Work
2.1 Character Understanding and its evolution with LLM
Character understanding has made significant contributions to advancing story comprehension in psychological, literary, and educational research  Bower and Morrow (1990); McKee (2010); Paris and Paris (2003). This multifaceted field encompasses a variety of tasks, including character recognition, coreference resolution, summarization, role detection, goal understanding, personality analysis, and question answering Sang et al. (2022b); Bamman et al. (2020); Chen et al. (2022); Frermann et al. (2018); Vallurupalli et al. (2024); Yu et al. (2023b). Recent studies have shown promising performance by LLMs on character understanding tasks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023). However, these studies often overlook the potential influence of memorization on model performance. Fictional works, including movie and TV show scripts, famous novels, and news texts, serve as primary sources for narrative understanding Sang et al. (2022a); Tapaswi et al. (2016); Caselli and Vossen (2017). These texts, much like copyrighted books Chang et al. (2023a), are prone to strong memorization. Naturally, this raises curiosity about how memorization affects downstream tasks, particularly character understanding. While previous research has demonstrated the memorization capabilities of LLMs, our work further explores its specific impact on character understanding and proposes an approach to mitigate this effect.
2.2 LLMs’ Memorization and Reasoning
Among LLMs’ various capabilities Li et al. (2024c, e); Zhang et al. (2024); Li et al. (2024d), the reasoning abilities Tong et al. (2023); Yang et al. (2023); Xiong et al. (2024b); Tong et al. (2024); Xiong et al. (2024a) of LLMs Li et al. (2024b); Tan et al. (2024) have garnered significant attention, but their memorization has also been shown to play a crucial role in influencing benchmark reasoning performance Zhou et al. (2023); Xie et al. (2024a); Chen et al. (2024); Li et al. (2024a); Biderman et al. (2024). The influence of memorization is complex: data contamination can inflate performance metrics Khandelwal et al. (2020); Blevins and Zettlemoyer (2022); Carlini et al. (2023), while knowledge conflict can reduce them Su et al. (2024); Wang et al. (2024); Xie et al. (2024b); Xu et al. (2024).
While prior work has focused on detecting and mitigating memorization effects in reasoning tasks such as numerical, commonsense, and logical reasoning Xie et al. (2024a); Sainz et al. (2023); Li et al. (2024f), character understanding tasks introduce additional challenges due to their diverse formats and varying memory demands. Unlike structured reasoning tasks where excessive memorization is clearly undesirable, character understanding often requires a nuanced balance between recall and inference.
2.3 Verbatim and Gist Memory in LLM Reasoning
Cognitive science distinguishes between verbatim memory, which retains exact details, and gist memory, which captures the essential meaning Brainerd and Reyna (2002); Reyna and Brainerd (1998). For example, verbatim memory would store a sentence like "The black-and-white cat sat on a soft, red mat in the living room" with all details intact, while gist memory would reduce it to "A cat was sitting indoors," focusing on the core idea while discarding specifics. Human reasoning primarily relies on gist memory, as the two types of memory function independently rather than hierarchically Brainerd and Reyna (1992).
In LLM research, memorization is often equated with verbatim memory, where models retrieve and reproduce exact strings from pretraining data Carlini et al. (2023). A common mitigation strategy involves curating test samples that the model has never seen before Xie et al. (2024a). However, in character understanding tasks, acquiring obscure or newly released datasets can be costly and difficult to control in terms of quality. Instead, our work proposes leveraging gist memory to reduce reliance on precise memorization, encouraging LLMs to generalize character understanding beyond string-matching. By shifting the focus toward gist memory, we aim to align LLM character analysis with human-like reasoning rather than dataset recall.
@@ 3 Data and Experiment Setup
In our work, we utilize two distinct datasets to evaluate LLMs on: (1) their reasoning capabilities in character understanding under gist memory and (2) their memorization of original source texts.
3.1 Benchmark Datasets for Character Understanding Tasks
To evaluate the interplay between memorization and reasoning in character understanding tasks, we employ six publicly available datasets covering diverse task types. These datasets span different narrative mediums, including TV shows, movies, and novels, ensuring diversity in both content and character representation. The task formulations and dataset compositions are detailed below, with an overview provided in Table 1. Further task details are available in  6.1.
Datasets
We evaluate character understanding using six publicly available datasets spanning diverse sources, including TV shows, movies, and novels. For coreference resolution, we use the latest release of the MovieCoref dataset  [^1] https://github.com/usc-sail/mica-character-coref, which provides annotated coreference chains in movie scripts. For character guess, we adopt TVSHOWGUESS  [^2] https://github.com/YisiSang/TVSHOWGUESS, where models must identify the speaker of a given dialogue. For personality understanding, we use PERSONET  [^3] https://github.com/Gorov/personet_acl23, a dataset derived from literary sources with personality trait annotations. For role detection, we employ CSI  [^4] https://github.com/EdinburghNLP/csi-corpus, which focuses on identifying key characters in crime narratives. For open-domain question answering, we evaluate models on FriendsQA  [^5] https://github.com/emorynlp/FriendsQA, where questions are answered by extracting spans from TV show dialogues. Lastly, for summarization, we use SummScreen  [^6] https://github.com/mingdachen/SummScreen, which pairs TV show transcripts with corresponding summaries. We follow the original dataset authors’ preprocessing, development, and test splits to ensure comparability. Further dataset details are provided in Appendix G.
@@ 3.1 Benchmark Datasets for Character Understanding Tasks
We followed the original evaluation metrics used in the respective tasks, shown in Table 1.
Task Name Task Type Data Source Metrics
@@
3.2 Fictional Works Memorization Benchmark Dataset
This dataset is designed to evaluate LLMs’ memorization of fictional scripts by assessing their ability to recall verbatim details from diverse narrative sources.
Benchmark Sources
The six benchmark datasets described in Section 6.1 are derived from a diverse set of sources, including 91 TV shows, 12 movies, and 33 classic literary works.
Dataset Construction
To ensure broad coverage across different narrative mediums, we sample 200 script segments from 40 selected sources, comprising 11 TV series, 9 movies, and 20 novels. Each source contributes 5 excerpts, with an average segment length of 312 words.
Segment Format
Each excerpt follows the format: TV Show Title; Scene, where the scene consists of a mixture of dialogue and narration extracted from the original script.
@@ 3.3 Models
we evaluate both open-source and closed-source models, leveraging their APIs for inference. Our evaluation includes three frontier LLMs: the open-source Llama 3.3-70B AI (2024) and DeepSeek V3 Liu et al. (2024), as well as the closed-source GPT-4o OpenAI (2023).
@@ 4 Soft Setting: Memorization-Controlled Prompting in Character Guessing
First, we investigate whether prompting strategies can subtly influence LLMs’ response patterns in a soft setting, where the model is not explicitly restricted from using memorization but is instead encouraged to rely more on reasoning over direct retrieval.
@@ 4.1 Character Guessing Task Format
Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where each utterance $U_{i}$ is paired with a speaker label $S_{i}$. In the Character Guessing Task, the model is given a dialogue excerpt without speaker annotations and must predict the correct speaker for each utterance, $f_{\text{guess}}(U_{i})\rightarrow S_{i}$. We follow the original format and setting Sang et al. (2022b), using the TVShow Guess Dataset described in Section 3.1. This task is designed to prompt the model to make guesses based on character relationships, key events, and personality traits within the text. However, its format closely resembles the name cloze tests Chang et al. (2023b) recently used to evaluate LLM memorization.
@@ 4.2 Prompting Strategies
We design two distinct prompting conditions to examine how LLMs rely on different memory mechanisms when identifying characters. Verbatim prompting encourages mechanical memorization, where the model recalls exact substrings from its training data and reconstructs missing parts without reasoning. In contrast, gist prompting focuses on extracting key information and performing memory-based matching. For character understanding tasks, the most relevant gist cues typically include character relationships, significant events, and personality traits.
@@
Verbatim Prompt: "…Do this by direct memory retrieval—do not infer based on traits, interactions, or context. Recall names exactly from past dialogue knowledge without explanations or reasoning."
Gist Prompt: "…Do this by analyzing relationships, key events, and personality traits to match them with known characters. Do not directly retrieve from memory who said the sentences."
@@ 4.3 Result and Analysis
Figure 2 visualizes the performance of the three models on the Character Guessing task under different prompting conditions. The performance trends remain consistent across models, and our key observations are as follows:
  • Baseline Performance: The baseline condition reflects the overall performance of the models, with GPT-4o demonstrating a clear lead, while Llama 3.3-70B and DeepSeek V3 exhibit similar performance.
@@
  • Effect of Gist Prompting: The Gist condition results in a non-trivial performance drop, suggesting that our prompt effectively reduces the models’ reliance on memorization when generating responses.

### v5

@@ Abstract
Large Language Models (LLMs) achieve striking performance on character understanding tasks, yet this success raises a key concern: are they truly reasoning about fictional characters, or merely recalling them from memorization? While prior work has explored memorization in general reasoning tasks, it remains unclear how memorization affects character understanding. To investigate this, we draw on cognitive science distinctions between "verbatim" (exact recall) and "gist" (semantic abstraction) memorization, and introduce a graded intervention framework with two levels of disruption. The soft setting preserves surface text but applies prompting strategies to encourage reasoning over recall, while the hard setting perturbs key character references to directly block memorization cues. Evaluating GPT-4o, LLaMA3.3-70B, and DeepSeek V3 on six character understanding benchmarks, we find that our hard setting leads to a dramatic performance drop of 45.2%—revealing substantial reliance on verbatim memorization and data leakage in existing benchmarks. In contrast, the soft setting steers models to reason more robustly without harming performance, showing that prompting alone can shift LLMs between memorization-based and reasoning-based strategies.
@@ 1 Introduction
Character understanding tasks assess a model’s ability to comprehend fictional characters’ personalities, goals, roles, and relationships within a narrative Yu et al. (2023a); Vallurupalli et al. (2024); Shen et al. (2024); Brahman et al. (2021). Recent LLMs have shown strong performance on such benchmarks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023), raising an important question: do these models truly understand characters, or are they simply recalling text seen during pretraining Chang et al. (2023a)? For instance, in a speaker guessing task from The Big Bang Theory (Figure 1), a model may identify Sheldon correctly not by reasoning over past actions (e.g., preparing lessons for Howard), but by matching exact phrases seen in previous data.
This concern is especially relevant in character understanding, as many benchmarks are based on widely consumed fictional works, such as TV shows, movies, or novels Sang et al. (2022a). Given the massive and opaque pretraining corpora of LLMs, models are likely to have encountered these works verbatim. As illustrated in Figure 1, such memorization can yield high performance even without genuine understanding. As a result, performance may reflect reliance on what cognitive psychology calls verbatim memorization—exact recall of surface text Huang et al. (2024); Ippolito et al. (2022)—rather than gist memorization, which captures semantic understanding and general patterns Reyna and Brainerd (1998). This challenges the validity of current evaluations: are models demonstrating true reasoning, or simply relying on verbatim memorization to regurgitate answers? In this work, we formulate a testable hypothesis: if LLMs’ strong performance is largely driven by memorized surface cues, then removing those cues—such as character names—should lead to a significant drop in accuracy.
Figure 1: Gist vs. verbatim memorization in speaker identification under original and anonymized scripts. This figure compares two memorizaiton strategies in identifying speakers (P0, P1) from a dialogue. On the left, character names are intact; on the right, all names are anonymized. The gist-based approach (blue) is prompted to "analyze personality traits, tones, and relationships", and infers speaker identity through reasoning—remaining robust even when names are masked. In contrast, the verbatim-based approach (yellow) is prompted to "recall names from exact dialogue memorization", and may misattribute the script (e.g., predicting Friends instead of The Big Bang Theory) when superficial cues are removed.
To investigate this hypothesis, we develop a graded intervention framework with two levels of disruption: a soft setting and a hard setting. The soft setting retains the original input scripts but uses gist-based prompting to guide the model toward abstraction and inference, without removing surface-level cues. In contrast, the hard setting replaces all character names with culturally distinct alternatives, aiming to block verbatim memorization while preserving the narrative content. We apply this graded intervention framework to six character understanding tasks and evaluate three LLMs: GPT-4o, LLaMA3-70B, and DeepSeek V3.
Our experiments provide clear support for our initial hypothesis: removing verbatim memorization cues significantly reduces LLMs’ performance on character understanding tasks, confirming that models heavily rely on memorized surface content. Specifically, we observe performance drops of up to 45.8% when verbatim memorization cues are blocked. In contrast, gist-based prompting maintains accuracy by steering model behavior toward reasoning rather than verbatim recall. These results suggest that current benchmarks may overestimate models’ genuine reasoning capabilities and that prompting strategies can effectively disentangle reasoning from memorization.
In summary, our key contributions are:
@@
    Identifying Memorization in Character Understanding: We empirically demonstrate significant performance drops (up to 46.2%) across benchmarks when verbatim memorization cues are disrupted, revealing a heavy reliance on surface-level recall rather than genuine reasoning.
@@
    A Graded Framework to Disentangle Reasoning from Memorization: Inspired by cognitive distinctions, we introduce a soft–hard intervention framework. The soft setting employs gist-based prompting to encourage abstraction, while the hard setting explicitly blocks verbatim recall through surface-level perturbation (Section 6).
@@
    Prompting as a Diagnostic Tool: We show that prompting style profoundly influences model behavior: verbatim prompts reinforce memorization, whereas gist prompts facilitate deeper inference, providing a practical approach to probe underlying reasoning processes in LLMs (Section 4.3).
@@ 2 Related Work
2.1 Character Understanding and Its Evaluation with LLMs
Character understanding has long been a focus in psychological, literary, and educational research Bower and Morrow (1990); McKee (2010); Paris and Paris (2003). It spans tasks such as character recognition, coreference resolution, summarization, role detection, goal modeling, personality analysis, and question answering Bamman et al. (2020); Chen et al. (2022); Frermann et al. (2018); Vallurupalli et al. (2024); Yu et al. (2023b).
Recent work demonstrates that LLMs achieve strong results on character-centric benchmarks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023). However, many of these datasets are derived from publicly available or widely consumed fictional sources, such as movie scripts, TV shows, and novels Sang et al. (2022a); Tapaswi et al. (2016); Caselli and Vossen (2017), which makes them vulnerable to pretraining overlap and surface-level memorization Chang et al. (2023a). Although prior research has studied LLM memorization broadly Carlini et al. (2023); Xie et al. (2024); Yang et al. (2023a), little is known about how it specifically affects character understanding tasks. Our work fills this gap by examining memorization’s role in these tasks and proposing interventions to distinguish between recall and reasoning.
2.2 Memorization and Reasoning in LLMs
LLMs’ impressive reasoning capabilities have attracted significant attention Tong et al. (2023); Yang et al. (2023b); Xiong et al. (2024); Tan et al. (2024); Jiang et al. (2025), yet their behavior is also strongly influenced by memorization Zhou et al. (2023); Chen et al. (2024); Biderman et al. (2024). Prior studies indicate that memorization of training data can artificially inflate benchmark results Khandelwal et al. (2020); Blevins and Zettlemoyer (2022), while conflicting memorized content can negatively affect performance Su et al. (2024); Wang et al. (2024); Yang et al. (2025). Unlike structured reasoning tasks, character understanding tasks inherently involve some memorization (e.g., narrative consistency), complicating the distinction between genuine reasoning and mere recall.
Cognitive science provides a useful conceptual framework for analyzing memory, distinguishing between verbatim memorization—the precise recall of surface-level details—and gist memorization—the abstraction of semantic meaning Reyna and Brainerd (1998); Brainerd and Reyna (2002). To avoid confusion with architectural notions of "memory" in NLP and machine learning Alizadeh et al. (2024); Kwon et al. (2023), we consistently adopt these terms to describe different recall styles rather than neural memory mechanisms. In our work, this distinction serves as the theoretical foundation for disentangling shallow memorization from deeper reasoning behavior in large language models.
Recent NLP research has formalized methods for quantifying and disentangling memorization from reasoning Lesci et al. (2024); Jin et al. (2024); Carlini et al. (2023); Prabhakar et al. (2024), examining phenomena such as multifaceted memorization Prashanth et al. (2024); Li et al. (2024) and logit-based tracing Chen et al. (2024). Mitigation strategies (e.g., out-of-distribution filtering Biderman et al. (2024)) are challenging to apply to character-centric tasks, which typically rely on known fictional content. Thus, we propose a prompting-based approach designed specifically to reduce verbatim reliance and promote gist-level reasoning in character understanding tasks.
@@ 3 Data and Experiment Setup
To comprehensively evaluate character understanding in LLMs, we compile two datasets: CharBench and CharScript. Inspired by NLP benchmark collections such as GLUE Wang et al. (2018), our CharBench integrates six publicly available datasets spanning character-centric tasks like coreference resolution, personality inference, and dialogue-based QA (details in Section 6.1).
The second dataset, CharScript, is newly constructed for this study by systematically sampling textual segments from original fictional sources underlying CharBench, explicitly probing models’ verbatim memorization (details in Section 3.2).
3.1 CharBench: A Benchmark Collection for Character Understanding
CharBench aggregates six widely recognized NLP benchmarks covering diverse narrative mediums—TV shows, movies, and novels. To ensure comparability, we retain each dataset’s original splits, preprocessing, and evaluation metrics. Detailed task descriptions and implementations are provided in Appendix E.
Benchmark Sources
Following established NLP benchmark practices Wang et al. (2018), CharBench includes:
  • MovieCoref Chen and Choi (2016): Coreference resolution in movie scripts.
  • TVSHOWGUESS Sang et al. (2022b): Speaker identification in TV dialogues.
  • PERSONET Yu et al. (2023b): Personality inference from literary passages.
  • CSI Corpus Frermann et al. (2018): Role detection from crime shows.
  • FriendsQA Yang and Choi (2019): Multi-turn dialogue-based QA.
  • SummScreen Chen et al. (2022): TV episode summarization.
@@ 3.1 CharBench: A Benchmark Collection for Character Understanding
We strictly follow original dataset metrics (exact match, ROUGE, F1 score), as detailed in Table 1.
Task Name Task Type Data Source Metric
@@
3.2 CharScript: Fictional Works Memorization Benchmark Dataset
Motivated by recent NLP research on memorization in language models Chang et al. (2023b), we constructed the CharScript dataset to evaluate whether LLMs can recognize the source of fictional text segments. Following a similar setup to Chang et al. (2023b), we probe whether models can correctly attribute short passages to their original narrative sources, as an indicator of verbatim memorization. We present our experimental design and findings in Section 5.1.
We systematically sampled and extracted 200 textual segments, averaging approximately 312 words each, from 40 diverse fictional sources (11 TV series, 9 movies, and 20 novels). These sources were carefully selected to maximize narrative diversity, textual complexity, and representativeness of popular fictional works included in CharBench. Each source contributes exactly 5 segments, structured uniformly in the format [Title]; Scene context, combining both original dialogue and narrative descriptions.
@@ 3.3 Models
We evaluate three frontier LLMs—open-source LLaMA 3.3-70B AI (2024) and DeepSeek V3 Liu et al. (2024), as well as the closed-source GPT-4o OpenAI (2023)—via API-based inference. To address stochastic variability, each model was run three times per setting, and average performance is reported. The total cost of our experiments was approximately \$100.
@@ 4 Soft Setting: Memorization-Controlled Prompting in Character Guessing
In the introduction, we formulated a testable hypothesis that LLMs’ strong performance on character understanding tasks relies significantly on verbatim memorization. As a first step toward testing this hypothesis, we investigate whether prompt design alone—without explicitly removing surface-level textual cues—can shift model responses from verbatim recall toward gist-based reasoning.
@@ 4.1 Character Guessing Task Format
We select the Character Guessing Task specifically because it closely mirrors the name cloze tests Chang et al. (2023b) previously established as effective probes for verbatim memorization in LLMs. Name cloze tests measure memorization by removing character names from familiar contexts and testing if models can accurately recall them. Similarly, in our task, speaker labels are removed from dialogue excerpts, requiring the model to infer speakers solely from textual context. If models rely heavily on verbatim memorization, accuracy should significantly degrade when surface cues—such as explicit character names—are disrupted. Conversely, if models genuinely reason about character interactions and relationships, performance should remain relatively stable despite these perturbations.
Following the original TVShow Guess setup Sang et al. (2022c), the model is tasked with predicting the speaker of each utterance in a dialogue segment where all speaker labels are removed.
@@ 4.2 Prompting Strategies
We design two distinct prompting conditions to examine how LLMs rely on different memorization mechanisms when identifying characters. Verbatim prompting encourages mechanical memorization, where the model recalls exact substrings from its training data and reconstructs missing parts without reasoning. In contrast, gist prompting focuses on extracting key information and performing memorization-based matching. For character understanding tasks, the most relevant gist cues typically include character relationships, significant events, and personality traits.
@@
Verbatim Prompt: "…Do this by direct memorization retrieval—do not infer based on traits, interactions, or context. Recall names exactly from past dialogue knowledge without explanations or reasoning."
Gist Prompt: "…Do this by analyzing relationships, key events, and personality traits to match them with known characters. Do not directly retrieve from memorization who said the sentences."
@@ 4.3 Result and Analysis
Our key observations are as follows:
  • Baseline Performance: The baseline condition reflects the overall performance of the models, with GPT-4o demonstrating a clear lead to others, while Llama 3.3-70B and DeepSeek V3 exhibit quite similar performance.
@@
  • Effect of Gist Prompting: The Gist condition results in a non-trivial performance drop, approximately 10%, suggesting that our prompt effectively reduces the model’s dependence on memorization when generating responses.
Our results indicate that gist-based prompting can indeed shift model responses toward reasoning, as evidenced by the performance drop under the Gist condition (see Figure 2). However, the moderate performance drop (approximately 10%) suggests that even gist prompting may not completely eliminate the models’ reliance on verbatim memorization. Previous research similarly indicates that prompting without explicit reasoning steps may fail to fully engage true reasoning processes Yu (2025). Therefore, to conclusively test our hypothesis—that models’ strong benchmark performance relies significantly on verbatim memorization—we introduce a more direct and explicit intervention aimed at disrupting memorized surface-level cues.